# Violation of Quantum Bilocal Inequalities on Mutually-Commuting von Neumann Algebra Models

### 01 Apr 2026 v2

@@ 4. Bilocal inequalities and their bounds
(4.1) $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}.$
Analogously, corresponding to Ineq. (3.2), we define
(4.2) $\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|,$
where, $K_{\tau}=\tau(A_{0}(B_{0}+B_{1})C_{0}),\ L_{\tau}=\tau(A_{1}(B_{0}-B_{1})C_{1})$.
We say that the state $\tau$ together with the observables $A_{x}$, $B_{y}$, $C_{z}$ satisfies the bilocal inequality if $\mathcal{S}_{\tau}\leq 2$, and violates it if $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$. The same criterion applies to $\mathcal{S}_{\tau}^{\prime}$ in Ineq. (4.2).
Next, we rely on the abelianness of the algebra to determine the bounds for the bilocal inequality, respectively.
The following conclusion indicates that in the mutually-commuting von Neumann algebra model, the supremum of $\mathcal{S}_{\tau}$ defined in Ineq. (4.1) is $2\sqrt{2}$. This coincides with the case in non-relativistic quantum mechanics, where the bilocal quantity $\mathcal{S}$ in Ineq. (3.1) attains a maximal violation of $2\sqrt{2}$ based on quantum resources.
Theorem 1. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
Proof. According to the Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{\mathcal{ABC}}\rightarrow\mathcal{B}(\mathcal{H}_{\tau})$ and a cyclic vector $\Omega\in\mathcal{B}(\mathcal{H}_{\tau})$ such that the set $\{\pi_{\tau}(O)\Omega:\ O\in\mathcal{M}_{\mathcal{ABC}}\}$ is dense in $\mathcal{H}_{\tau}$. It follows by applying the Cauchy-Schwarz inequality that
@@
[…]
Building on the results above, we now investigate how the quantity $\mathcal{S}_{\tau}$ in Ineq. (4.1) depends on the abelianness of the algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, and $\mathcal{M}_{\mathcal{C}}$. Specifically, the results indicate that in entanglement swapping networks, the abelianness of the three algebras plays distinct roles in reducing the upper bound of the inequality to 2, i.e., determining the conditions under which no violation of the bilocal inequality can occur. This is not a simple generalization of the bipartite Bell scenario [27], where, with only two systems, Summers et al. showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
Theorem 2. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
(1) if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2;$
(2) if $\mathcal{M}_{\mathcal{B}}$ is Abelian, then
$\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|\leq 2.$
Proof. To see (1), since $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, respectively, the eight elements
@@
where the first inequality follows the fact that $I\leq B_{i}\leq I\ (i=0,1)$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy Schwarz inequality and the non-negativeness of $A_{\epsilon_{0}\epsilon_{1}}\ \text{and}\ \ B_{\epsilon_{0}\epsilon_{1}}$.
To prove (2), since $\mathcal{M}_{\mathcal{B}}$ is Abelian, the four elements
$B_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}a_{0})(1+\epsilon_{1}a_{1}),$
with $\epsilon_{0},\epsilon_{1}\in\{+,-\}$ are positive. It follows from a direct computation that
$B_{0}+B_{1}=2(B_{++}-B_{--}),\ B_{0}-B_{1}=2(B_{+-}-B_{-+}).$
So one obtains that
$\mathcal{S}_{\tau}^{\prime}$ $=$ $\tau[A_{0}(B_{0}+B_{1})C_{0}]+\tau[A_{1}(B_{0}-B_{1})C_{1}]$
$=$ $2(\tau[A_{0}(B_{++}-B_{--})C_{0}]+\tau[A_{1}(B_{+-}-B_{-+})C_{1}])$
$\leq$ $2(\tau[B_{++}]+\tau[B_{--}]+\tau[B_{+-}]+\tau[B_{-+}])=2$
[…]
Remark. The above theorem illustrates a phenomenon: the violations of the two Bell-like inequalities, i.e., $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$ and $2<\mathcal{S}^{\prime}_{\tau}\leq 2\sqrt{2}$ can serve as an indicator of the non-abelianness of the underlying algebras. Applying the theorem, we can infer from a violation of Ineq. (4.1) that at least one of the algebras $\mathcal{M_{A}}$ and $\mathcal{M_{C}}$ is non-abelian; Similarly, a violation of Ineq. (4.2) implies that $\mathcal{M_{B}}$ is non-abelian. Here, one also needs to consider the chosen observables and the specific state.
Nevertheless, it is important to note that these conclusions depend not only on the algebraic structure but also on the specific choice of observables and the particular state under consideration. To isolate the influence of the algebraic structure and examine separately the role of observables and the state in determining the values $\mathcal{S}_{\tau}$ and $\mathcal{S}_{\tau}^{\prime}$ in Ineqs. (4.1) and  (4.2), the following conclusion can be regarded as a generalization of the results in the non-relativistic quantum mechanical framework. For the sake of clarity, we introduce a tensor product structure into the TMCvNA model. Before presenting the following result, we first introduce the definition of separable states.
Definition 3. (Separable states) Let $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ be mutually commuting von Neumann subalgebras of $\mathcal{M}_{\mathcal{AB}}=(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}$. A state $\tau\in\mathcal{M}_{\mathcal{AB}}^{*}$ is said to be separable if there exist states $\xi_{\alpha}\in\mathcal{M}_{\mathcal{A}}^{*}$, $\eta_{\alpha}\in\mathcal{M}_{\mathcal{B}}^{*}$, and positive coefficients $\lambda_{\alpha}$ with $\sum_{\alpha}\lambda_{\alpha}=1$ such that for all $A\in\mathcal{M}_{\mathcal{A}},B\in\mathcal{M}_{\mathcal{B}}$,
$\tau(AB)=\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A)\eta_{\alpha}(B).$
[…]
$\sup_{\{A_{x},B_{y},C_{z}\}}\mathcal{S}_{\tau}=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})=2.$
[…]
$=\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A_{0}+A_{1})\eta_{\alpha}(B_{0}^{L})||\sum_{\beta}\mu_{\beta}M_{\beta}(B_{0}^{R})N_{\beta}(C_{0}+C_{1})|}$
$\hskip 18.49988pt+\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A_{0}-A_{1})\eta_{\alpha}(B_{1}^{R})||\sum_{\beta}\mu_{\beta}M_{\beta}(B_{1}^{R})N_{\beta}(C_{0}-C_{1})|}$
$\leq\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(A_{0}+A_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(C_{0}+C_{1})|)}+\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(A_{0}-A_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(C_{0}-C_{1})|)}$
$\leq\sqrt{\sum_{\alpha}\lambda_{\alpha}(|\xi_{\alpha}(A_{0}+A_{1})|+|\xi_{\alpha}(A_{0}-A_{1})|)}\sqrt{\sum_{\beta}\mu_{\beta}(|N_{\beta}(C_{0}+C_{1})|+|N_{\beta}(C_{0}-C_{1})|)}$
$=2\sqrt{\sum_{\alpha}\lambda_{\alpha}\max\{\xi_{\alpha}(A_{0}),\xi_{\alpha}(A_{1})\}}\sqrt{\sum_{\beta}\mu_{\beta}\max\{N_{\beta}(C_{0}),N_{\beta}(C_{1})\}}$
$\leq 2.$
[…]
From the above conclusion, we see that whether a bilocal inequality can be violated depends on the choice of observables and the state, and even more so on the structure of the algebra. In the following, we define a quantity
$\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})$
that appears in Theorem 3.
Combining Theorems 1, 2 and 3, one naturally obtains the following corollary. A similar conclusion can be drawn for the quantity $\mathcal{S}^{\prime}_{\tau}$ in Ineq. (4.2).
Corollary 4. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
@@
where $k$ is a positive constant. Consequently, the functional $\phi\rightarrow\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
Proof. To show (1), note that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and combining this with the proof of Theorem 2, we obtain (1).
(2) holds by the fact that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and by Theorem 2.
@@
[…]
5. Maximal violation of bilocal inequalities and algebraic structures
In this section, we aim to point out that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra. Here, violation refers to exceeding the upper bound of 2 for the $\mathcal{S}_{\tau}$ in inequality (4.1), while maximal violation means attaining the value $2\sqrt{2}$ for the same quantity.
In the following theorem, we attempt to analyze the conditions for maximal violation.
Theorem 5. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if $\tau$ is faithful, then
@@ 5. Maximal violation of bilocal inequalities and algebraic structures
Remark. This theorem states that when the inequality is maximally violated, $A_{0},\ A_{1}$ and $\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices in $\mathcal{M}_{\mathcal{A}}$ and generate a copy of $M_{2}(\mathbb{C})$ in $\mathcal{M}_{\mathcal{A}}$. The same holds for $C_{0},\ C_{1}$ and $\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{M}_{\mathcal{C}}$. Moreover, $\tau(A_{0}A_{1}+A_{1}A_{0})=0,\ \tau(C_{0}C_{1}+C_{1}C_{0})=0.$
Proof. If $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$, it follows from the proof of Theorem 1 that for any $t\in[0,1]$, these equalities hold:
(5.1) $|\tau(B_{0}(A_{0}+A_{1})(C_{0}+C_{1}))|=|\tau(B_{1}(A_{0}-A_{1})(C_{0}-C_{1}))|,$
(5.2) $\pi_{\tau}(B_{0})\Omega=k_{0}\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega,$
(5.3) $\pi_{\tau}(B_{1})\Omega=k_{1}\pi_{\tau}[(A_{0}-A_{1})(C_{0}-C_{1})]\Omega,$
(5.4) $\|\pi_{\tau}(B_{0})\Omega\|=\|\pi_{\tau}(B_{1})\Omega\|=1,$
(5.5) $\tau[(A_{0}+A_{1})^{2}]=t\tau[(C_{0}+C_{1})^{2}],$
(5.6) $\tau[(A_{0}-A_{1})^{2}]=t\tau[(C_{0}-C_{1})^{2}],$
(5.7) $A_{0}^{2}+A_{1}^{2}=2I,C_{0}^{2}+C_{1}^{2}=2I.$

### 15 Apr 2026 v5

@@ 3 Bilocal inequalities and their bounds
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}.$ (3.1)
We say that the state $\tau$ together with the observables $A_{x}$, $B_{y}$, $C_{z}$ satisfies the bilocal inequality if $\mathcal{S}_{\tau}\leq 2$, and violates it if $\mathcal{S}_{\tau}>2$.
The following conclusion indicates that in the TMCvNA of entanglement swapping networks, the supremum of $\mathcal{S}_{\tau}$ defined in Eq. (3.1) is $2\sqrt{2}$. This coincides with the case in non-relativistic quantum mechanics, where the bilocal quantity $\mathcal{S}$ in Ineq. (2.1) attains a maximal violation of $2\sqrt{2}$ allowed by quantum resources.
Theorem 3.1.
For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, we always have $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
Proof According to the Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{\mathcal{ABC}}\rightarrow\mathcal{B}(\mathcal{H}_{\tau})$ and a cyclic vector $\Omega\in\mathcal{H}_{\tau}$ such that the set $\{\pi_{\tau}(O)\Omega:\ O\in\mathcal{M}_{\mathcal{ABC}}\}$ is dense in $\mathcal{H}_{\tau}$. It follows by applying the Cauchy-Schwarz inequality that
@@
The final inequality invokes the condition that $I\leq A_{i}\leq I,\ -I\leq C_{i}\leq I$ and the positivity property of $\tau$. $\Box$
Building on the results above, we now investigate how the quantity $\mathcal{S}_{\tau}$ in Eq. (3.1) depends on the abelianness of the algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, and $\mathcal{M}_{\mathcal{C}}$. Specifically, the results indicate that in entanglement swapping networks, the abelianness of the three algebras plays distinct roles in reducing the upper bound of the inequality to 2, i.e., determining the conditions under which no violation of the bilocal inequality can occur. This is not a simple generalization of the bipartite Bell scenario [47], where, with only two systems, Summers et al. showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
Theorem 3.2.
For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2.$
Proof Since $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, respectively, the eight elements
@@
where the first inequality follows the fact that $I\leq B_{i}\leq I\ (i=0,1)$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy-Schwarz inequality and the non-negativeness of $A_{\epsilon_{0}\epsilon_{1}}\ \text{and}\ \ C_{\epsilon_{0}\epsilon_{1}}$. $\Box$
The above theorem illustrates a phenomenon: the violation of the Bell-type inequality, i.e., $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$ can serve as an indicator of the non-abelianness of the underlying algebras. Applying the theorem, we can infer from a violation of the inequality in Theorem 3.2 that at least one of the algebras $\mathcal{M_{A}}$ and $\mathcal{M_{C}}$ is non-abelian.
In the following, we define a quantity
$\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}).$
Combining Theorems 3.1 and 3.2, one naturally obtains the following corollary.
Corollary 3.3.
In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
@@
where $k$ is a positive constant. Consequently, the functional $\phi\mapsto\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
Proof To show (1), note that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and combining this with the proof of Theorem 3.1, we obtain (1).
(2) holds by the fact that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and by Theorem 3.2.
@@
where the fourth one follows the Cauchy Schwarz inequality, and the last obeys the norm for elements of $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ are bounded. So $\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous in the state $\phi$. $\Box$
4 Maximal violation of bilocal inequalities and algebraic structures
In this section, we aim to show that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra. Here, violation refers to exceeding the bound for $\mathcal{S}_{\tau}$ in Eq. (3.1), while maximal violation means attaining the value $2\sqrt{2}$ for the same quantity.
The following theorem analyzes the conditions for maximal violation.
Theorem 4.1.
For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if the network state $\tau\in\mathcal{M}_{\mathcal{ABC}}^{*}$ is faithful, then
@@ 4 Maximal violation of bilocal inequalities and algebraic structures
Proof If $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$, it follows from the proof of Theorem 3.1 that for any $t\in[0,1]$, these equalities hold:
$|\tau(B_{0}(A_{0}+A_{1})(C_{0}+C_{1}))|=|\tau(B_{1}(A_{0}-A_{1})(C_{0}-C_{1}))|,$ (4.1)
$\pi_{\tau}(B_{0})\Omega=k_{0}\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega,$ (4.2)
$\pi_{\tau}(B_{1})\Omega=k_{1}\pi_{\tau}[(A_{0}-A_{1})(C_{0}-C_{1})]\Omega,$ (4.3)
$\|\pi_{\tau}(B_{0})\Omega\|=\|\pi_{\tau}(B_{1})\Omega\|=1,$ (4.4)
$\tau[(A_{0}+A_{1})^{2}]=t\tau[(C_{0}+C_{1})^{2}],$ (4.5)
$\tau[(A_{0}-A_{1})^{2}]=t\tau[(C_{0}-C_{1})^{2}],$ (4.6)
$A_{0}^{2}+A_{1}^{2}=2I,C_{0}^{2}+C_{1}^{2}=2I.$ (4.7)